OpenStudy (anonymous):

What is the factorization of the polynomial below? x2 + 11x + 10

OpenStudy (kropot72):

What two numbers multiplied together to get 10 add up to 11?

OpenStudy (anonymous):

We know it should be of the form \((x+a)(x+b)=x^2+(a+b)x+ab \equiv x^2 + 11x + 10\). We want \(a+b=11\) and \(ab=10\). What two numbers \(a\) and \(b\) can you think of that satisfy this?

jhonyy9 (jhonyy9):

(x+1)(x+10)=0